# ViWikiFC: Fact-Checking for Vietnamese Wikipedia-Based Textual Knowledge Source

### v1

@@ 5 Conclusions and Future Directions
In this paper, we construct the first Wikipedia Vietnamese corpus, open-domain and high-quality, including 20,916 samples to evaluate fact-checking models. Furthermore, we have discussed the data collection and annotation method and shared the insight we obtained during the annotation process, which can be applied in the development of non-English corpus creation. The best baseline and pipeline model (InfoXLM and BM25) only achieved 67.00% strict accuracy, making our feasible challenge for the Vietnamese language model in fact-checking tasks. We believe that ViWikiFC will encourage the development of Vietnamese fact-checking research.
@@
Following the development of fact-checking in English, we aim to develop our research for Vietnamese fact-checking by expanding our corpus from quantity to quality with more trustworthy data sources and constructing datasets not only on textual but also on images and tables for developing multi-modal fact-checking. Besides, we want our work to adapt to more advanced NLP tasks in low-resource language, such as fake news detection or machine reading comprehension, with advanced reasoning design and sub-task functionalities.
@@ Acknowledgement
This research was supported by The VNUHCM-University of Information Technology’s Scientific Research Support Fund.
@@
The corpora generated during and/or analyzed during the current study are available from the corresponding author on reasonable request.
@@ Appendix A Data Generation Rules Analysis
Table 13: SUPPORTS rules and examples for creating evidence (E) - claim (C) pairs. Simply, we only mention one rule to be applied in each example.
Rule Example Ratio
@@
Turn nouns into relative clauses E: Trong bốn nước xưa có nền văn minh lớn thì có ba nước xưa ở vào châu Á (´n Độ, Iraq (Lưỡng Hà) và Trung Quốc). (Among the four ancient civilizations, three were located in Asia (India, Iraq (Mesopotamia), and China).) C: Châu Á là nơi tồn tại ba nước có nền văn minh lớn. (Asia is the region where three major ancient civilizations existed.) 10.20%
Turn the object into relative clauses E: Lý thuyết ”Quả cầu tuyết Trái Đất” cho rằng những sự thay đổi về mức độ CO2 vừa là nguyên nhân gây ra, vừa là nguyên nhân làm kết thúc thời kỳ cực lạnh ở cuối Liên đại Nguyên Sinh (Proterozoic). (The ”Snowball Earth” theory posits that changes in the levels of CO2 were both the cause and the termination of the extreme cold period at the end of the Proterozoic Eon.) C: Theo lý thuyết ”Quả cầu tuyết Trái Đất” cho thấy rằng sự thay đổi thành phần không khí mà cụ thể là lưu lượng CO2 là tác nhân cho kỷ băng giá, và cũng kết thúc luôn thời kỳ cực lạnh ở kỷ Proterozoic. (The ”Snowball Earth” theory indicates that changes in the composition of the atmosphere, specifically the levels of CO2 which are the driving factors behind ice ages and the termination of extreme cold periods in the Proterozoic Eon.) 2.20%
@@
Rule Example Ratio
Use negative words (no, not, never, nothing, hardly, etc.) E: Sự kiện này dẫn tới việc Hiệp định Genève (1954) được ký kết và Việt Nam bị chia cắt thành hai vùng tập kết quân sự, lấy ranh giới là vĩ tuyến 17. (This event led to the signing of the Geneva Accords in 1954, which divided Vietnam into two military zones along the 17th parallel.) C: Sự kiện này dẫn tới việc Hiệp định Genève (1954) không được ký kết. (This event led to the Geneva Accords (1954) not being signed.) 23.00%
Replace words with antonyms E:Đội tuyển bóng nước Singapore đã giành huy chương vàng SEA Games lần thứ 27 vào năm 2017, tiếp tục chuỗi vô địch dài nhất của thể thao Singapore về môn môn cụ thể. (The Singapore water polo team won the gold medal at the 27th SEA Games in 2017, continuing Singapore’s longest winning streak in a specific sport.) C:Singapore liên tục gặp thất bại ở bộ môn bóng nước tại các kỳ SEA Games. (Singapore has continuously faced defeats in the sport of water polo at various SEA Games.) 14.00%
Opposite of quantity E: Năm 2010, tổng chi tiêu của Nhà nước vào khoa học và công nghệ chiếm khoảng 0.45% GDP. (In 2010, the government’s total expenditure on science and technology accounted for approximately 0.45% of the GDP.) C: Nhà nước chi ra hơn 2.00% GDP năm 2010 cho khoa học và công nghệ. (The government allocated over 2.00% of the GDP in 2010 for science and technology.) 9.80%
Opposite of time E: Năm 1860, dân số Singapore đã vượt quá 80,000 và hơn một nửa là người Hoa. (In 1860, the population of Singapore had exceeded 80,000, with over half being of Chinese descent.) C: Mãi đến năm 2000, Dân số Singapore mới đạt mốc 80,000. (It wasn’t until the year 2000 that Singapore’s population reached the 80,000 milestone.) 5.60%
Create a sentence that has the opposite meaning of a presupposition E: Đức đã tung ra 70.00% binh lực với các sư đoàn mạnh và tinh nhuệ nhất, chưa kể binh lực góp thêm của các nước đồng minh của Đức (Ý, Rumani, Bulgari, Hungary, Phần Lan…). (Germany deployed approximately 70.00% of its military power, not counting the additional military contributions from Germany’s allies (Italy, Romania, Bulgaria, Hungary, Finland…).) C: Đức không có bất kỳ đồng minh nào giúp sức. (Germany did not have any allies providing assistance.) 5.60%
Wrong reasoning about an object (House, car, river, sea, person, etc.) E: Khi gió mùa đổi hướng, các đường bờ biển giáp với biển Ả Rập và vịnh Bengal có thể phải hứng chịu xoáy thuận. (When the monsoon winds change direction, coastal areas bordering the Arabian Sea and the Bay of Bengal may be susceptible to cyclones.) C: Các đường bờ biển tiếp giáp ´n Độ có thể phải hứng chịu xoáy thuận khi gió mùa đổi hướng. (The coastal areas bordering India may be susceptible to cyclones when the monsoon winds change direction.) 23.60%
Wrong reasoning about an event E: Một trận động đất ở Valdivia, Chile với cường độ 9,4-9,6 độ richter, mức cao nhất từng được ghi nhận, khiến 1.000 đến 6.000 người chết. (An earthquake in Valdivia, Chile, with a magnitude of 9.4 to 9.6 on the Richter scale, the highest ever recorded, resulted in the deaths of approximately 1,000 to 6,000 people.) C: Cơn địa chấn ở Valvidia chỉ gây ra thiệt hại về tài sản. (The earthquake in Valdivia only caused property damage.) 30.80%
Others E: Tiếng Pháp không phải là ngôn ngữ chính thức ở Ontario, nhưng Đạo luật Dịch vụ Ngôn ngữ Pháp đảm bảo rằng các dịch vụ của tỉnh bang sẽ được cung cấp bằng ngôn ngữ này. (French is not the official language in Ontario, but the French Language Services Act ensures that provincial services will be provided in this language.) C: tiếng Pháp có là ngôn ngữ chung của bang Ontario, thì nó cũng sẽ bị loại bỏ khỏi hệ thống các dịch vụ. (Even though French is one of the official languages of the province of Ontario, it will be removed from the system of services.) 0.20%

### v2

@@ 5.  Conclusions and Future Directions
In this paper, we construct the first open-domain and high-quality Wikipedia Vietnamese corpus, including 20,916 samples to evaluate fact-checking models. Furthermore, we have discussed the data collection and annotation method and shared the insight we obtained during the annotation process, which can be applied in the development of non-English corpus creation. The best baseline and pipeline model (InfoXLM and BM25) only achieved 67.00% strict accuracy, making it a feasible challenge for the Vietnamese language model in fact-checking tasks. We believe that ViWikiFC will encourage the development of Vietnamese fact-checking research.
@@
Following the development of fact-checking in English, we aim to develop our research for Vietnamese fact-checking by expanding our corpus from quantity to quality with more trustworthy data sources and constructing datasets not only on textual but also on images and tables for developing multi-modal fact-checking. Besides, we want our work to adapt to more advanced NLP tasks in low-resource languages, such as fake news detection or machine reading comprehension, with advanced reasoning design and sub-task functionalities.
@@ Acknowledgement
This research is funded by Vietnam National University Ho Chi Minh City (VNU-HCM) under the grant number DS2025-26-01.
@@
Data available on reasonable request.
@@
6.  Bibliographical References
  • R. Aly, Z. Guo, M. S. Schlichtkrull, J. Thorne, A. Vlachos, C. Christodoulopoulos, O. Cocarascu, and A. Mittal (2021) FEVEROUS: fact extraction and verification over unstructured and structured information. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 1), Cited by: §2.1.
  • I. Augenstein, C. Lioma, D. Wang, L. Chaves Lima, C. Hansen, C. Hansen, and J. G. Simonsen (2019) MultiFC: a real-world multi-domain dataset for evidence-based fact checking of claims. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Hong Kong, China, pp. 4685–4697. External Links: Link, Document Cited by: §2.1.
  • W. Chang, X. Y. Felix, Y. Chang, Y. Yang, and S. Kumar (2019) Pre-training tasks for embedding-based large-scale retrieval. In International Conference on Learning Representations, Cited by: §2.2.
  • D. Chen, A. Fisch, J. Weston, and A. Bordes (2017) Reading Wikipedia to answer open-domain questions. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), R. Barzilay and M. Kan (Eds.), Vancouver, Canada, pp. 1870–1879. External Links: Link, Document Cited by: §2.2.
  • W. Chen, H. Wang, J. Chen, Y. Zhang, H. Wang, S. Li, X. Zhou, and W. Y. Wang (2020) TabFact: a large-scale dataset for table-based fact verification. In International Conference on Learning Representations, External Links: Link Cited by: §2.1.
  • Z. Chi, L. Dong, F. Wei, N. Yang, S. Singhal, W. Wang, X. Song, X. Mao, H. Huang, and M. Zhou (2021) InfoXLM: an information-theoretic framework for cross-lingual language model pre-training. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, K. Toutanova, A. Rumshisky, L. Zettlemoyer, D. Hakkani-Tur, I. Beltagy, S. Bethard, R. Cotterell, T. Chakraborty, and Y. Zhou (Eds.), Online, pp. 3576–3588. External Links: Link, Document Cited by: §1.
  • A. Conneau, K. Khandelwal, N. Goyal, V. Chaudhary, G. Wenzek, F. Guzmán, E. Grave, M. Ott, L. Zettlemoyer, and V. Stoyanov (2020) Unsupervised cross-lingual representation learning at scale. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, Online, pp. 8440–8451. External Links: Link, Document Cited by: §1, §2.2.
  • J. Devlin, M. Chang, K. Lee, and K. Toutanova (2019) BERT: pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), Minneapolis, Minnesota, pp. 4171–4186. External Links: Link, Document Cited by: §2.2.
  • J. L. Elman (1990) Finding structure in time. Cognitive Science 14 (2), pp. 179–211. External Links: ISSN 0364-0213, Document, Link Cited by: §2.2.
  • J. L. Fleiss (1971) Measuring nominal scale agreement among many raters.. Psychological Bulletin 76, pp. 378–382. External Links: Link Cited by: §3.1.
  • A. Graves and J. Schmidhuber (2005) Framewise phoneme classification with bidirectional lstm and other neural network architectures. Neural networks 18 (5-6), pp. 602–610. Cited by: §1, §2.2.
  • Z. Guo, M. Schlichtkrull, and A. Vlachos (2022) A survey on automated fact-checking. Transactions of the Association for Computational Linguistics 10, pp. 178–206. External Links: Link, Document Cited by: §1.
  • A. Gupta and V. Srikumar (2021) X-fact: a new benchmark dataset for multilingual fact checking. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 2: Short Papers), Online, pp. 675–682. External Links: Link, Document Cited by: §2.1.
  • V. Gupta, M. Mehta, P. Nokhiz, and V. Srikumar (2020) INFOTABS: inference on tables as semi-structured data. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, Online, pp. 2309–2324. External Links: Link, Document Cited by: §2.1.
  • A. Hanselowski, C. Stab, C. Schulz, Z. Li, and I. Gurevych (2019) A richly annotated corpus for different tasks in automated fact-checking. In Proceedings of the 23rd Conference on Computational Natural Language Learning (CoNLL), Hong Kong, China, pp. 493–503. External Links: Link, Document Cited by: §2.1.
  • B. Horne, S. Khedr, and S. Adali (2018) Sampling the news producers: a large news and feature data set for the study of the complex media landscape. In Proceedings of the International AAAI Conference on Web and Social Media, Vol. 12. Cited by: §1.
  • H. Hu, K. Richardson, L. Xu, L. Li, S. Kuebler, and L. Moss (2020) OCNLI: original chinese natural language inference. In Findings of EMNLP, External Links: Link Cited by: §3.3.
  • X. Hu, Z. Guo, G. Wu, A. Liu, L. Wen, and P. Yu (2022) CHEF: a pilot Chinese dataset for evidence-based fact-checking. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Seattle, United States, pp. 3362–3376. External Links: Link, Document Cited by: §2.1, §2.2.
  • T. V. Huynh, K. V. Nguyen, and N. L. Nguyen (2022) ViNLI: a Vietnamese corpus for studies on open-domain natural language inference. In Proceedings of the 29th International Conference on Computational Linguistics, Gyeongju, Republic of Korea, pp. 3858–3872. External Links: Link Cited by: §1, §3.5.4.
  • Y. Jiang, S. Bordia, Z. Zhong, C. Dognin, M. Singh, and M. Bansal (2020) HoVer: a dataset for many-hop fact extraction and claim verification. In Findings of the Association for Computational Linguistics: EMNLP 2020, Online, pp. 3441–3460. External Links: Link, Document Cited by: §2.1.
  • V. Karpukhin, B. Oguz, S. Min, P. Lewis, L. Wu, S. Edunov, D. Chen, and W. Yih (2020) Dense passage retrieval for open-domain question answering. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), B. Webber, T. Cohn, Y. He, and Y. Liu (Eds.), Online, pp. 6769–6781. External Links: Link, Document Cited by: §2.2.
  • J. D. M. C. Kenton and L. K. Toutanova (2019) BERT: pre-training of deep bidirectional transformers for language understanding. In Proceedings of NAACL-HLT, pp. 4171–4186. Cited by: §1.
  • J. Khouja (2020) Stance prediction and claim verification: an Arabic perspective. In Proceedings of the Third Workshop on Fact Extraction and VERification (FEVER), Online, pp. 8–17. External Links: Link, Document Cited by: §2.1.
  • Z. Lan, M. Chen, S. Goodman, K. Gimpel, P. Sharma, and R. Soricut (2019) ALBERT: a lite bert for self-supervised learning of language representations. In International Conference on Learning Representations, Cited by: §2.2.
  • K. Lee, M. Chang, and K. Toutanova (2019) Latent retrieval for weakly supervised open domain question answering. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, Florence, Italy, pp. 6086–6096. External Links: Link, Document Cited by: §2.2.
  • P. Lewis, E. Perez, A. Piktus, F. Petroni, V. Karpukhin, N. Goyal, H. Küttler, M. Lewis, W. Yih, T. Rocktäschel, et al. (2020) Retrieval-augmented generation for knowledge-intensive nlp tasks. Advances in Neural Information Processing Systems 33, pp. 9459–9474. Cited by: §1.
  • Y. Liu, M. Ott, N. Goyal, J. Du, M. Joshi, D. Chen, O. Levy, M. Lewis, L. Zettlemoyer, and V. Stoyanov (2019) RoBERTa: A robustly optimized BERT pretraining approach. CoRR abs/1907.11692. External Links: Link, 1907.11692 Cited by: §2.2.
  • R. Mahendra, A. F. Aji, S. Louvan, F. Rahman, and C. Vania (2021) IndoNLI: a natural language inference dataset for Indonesian. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, Online and Punta Cana, Dominican Republic, pp. 10511–10527. External Links: Link Cited by: §3.3.
  • J. Maillard, C. Gao, E. Kalbassi, K. R. Sadagopan, V. Goswami, P. Koehn, A. Fan, and F. Guzman (2023) Small data, big impact: leveraging minimal data for effective machine translation. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Toronto, Canada, pp. 2740–2756. External Links: Link, Document Cited by: §1.
  • T. Mikolov, I. Sutskever, K. Chen, G. S. Corrado, and J. Dean (2013) Distributed representations of words and phrases and their compositionality. Advances in neural information processing systems 26. Cited by: §1, §2.2.
  • D. Q. Nguyen and A. Tuan Nguyen (2020) PhoBERT: pre-trained language models for Vietnamese. In Findings of the Association for Computational Linguistics: EMNLP 2020, Online, pp. 1037–1042. External Links: Link, Document Cited by: §1, §2.2.
  • L. T. Nguyen and D. Q. Nguyen (2021) PhoNLP: A joint multi-task learning model for Vietnamese part-of-speech tagging, named entity recognition and dependency parsing. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Demonstrations, pp. 1–7. Cited by: §3.5.3.
  • R. Nogueira and K. Cho (2019) Passage re-ranking with bert. arXiv preprint arXiv:1901.04085. Cited by: §2.2.
  • R. Nogueira, W. Yang, K. Cho, and J. Lin (2019) Multi-stage document ranking with bert. arXiv preprint arXiv:1910.14424. Cited by: §2.2.
  • J. Nørregaard and L. Derczynski (2021) DanFEVER: claim verification dataset for Danish. In Proceedings of the 23rd Nordic Conference on Computational Linguistics (NoDaLiDa), Reykjavik, Iceland (Online), pp. 422–428. External Links: Link Cited by: §2.1.
  • L. Pan, X. Wu, X. Lu, A. T. Luu, W. Y. Wang, M. Kan, and P. Nakov (2023) Fact-checking complex claims with program-guided reasoning. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Toronto, Canada, pp. 6981–7004. External Links: Link, Document Cited by: §1, §4.2.3.
  • L. Phan, H. Tran, H. Nguyen, and T. H. Trinh (2022a) ViT5: pretrained text-to-text transformer for Vietnamese language generation. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies: Student Research Workshop, D. Ippolito, L. H. Li, M. L. Pacheco, D. Chen, and N. Xue (Eds.), Hybrid: Seattle, Washington + Online, pp. 136–142. External Links: Link, Document Cited by: §2.2.
  • Q. L. Phan, T. H. P. Doan, N. H. Le, N. B. D. Tran, and T. N. Huynh (2022b) Vietnamese sentence paraphrase identification using sentence-bert and phobert. In Intelligence of Things: Technologies and Applications, N. Nguyen, N. Dao, Q. Pham, and H. A. Le (Eds.), Cham, pp. 416–423. External Links: ISBN 978-3-031-15063-0 Cited by: §1.
  • C. Samarinas, W. Hsu, and M. L. Lee (2020) Latent retrieval for large-scale fact-checking and question answering with nli training. In 2020 IEEE 32nd International Conference on Tools with Artificial Intelligence (ICTAI), pp. 941–948. Cited by: §2.2.
  • C. Samarinas, W. Hsu, and M. L. Lee (2021) Improving evidence retrieval for automated explainable fact-checking. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies: Demonstrations, Online, pp. 84–91. External Links: Link, Document Cited by: §2.2.
  • T. Schuster, A. Fisch, and R. Barzilay (2021) Get your vitamin C! robust fact verification with contrastive evidence. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Online, pp. 624–643. External Links: Link, Document Cited by: §1, §2.1.
  • G. K. Shahi and D. Nandini (2020) FakeCovid–a multilingual cross-domain fact check news dataset for covid-19. arXiv preprint arXiv:2006.11343. Cited by: §1.
  • N. Thakur, N. Reimers, A. Rücklé, A. Srivastava, and I. Gurevych (2021) BEIR: a heterogeneous benchmark for zero-shot evaluation of information retrieval models. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 2), Cited by: §4.2.1.
  • J. Thorne, A. Vlachos, C. Christodoulopoulos, and A. Mittal (2018) FEVER: a large-scale dataset for fact extraction and VERification. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long Papers), New Orleans, Louisiana, pp. 809–819. External Links: Link, Document Cited by: §1, §1, §2.1, §2.2, §3.3, §3.3.
  • C. D. Tran, N. H. Pham, A. T. Nguyen, T. S. Hy, and T. Vu (2023) ViDeBERTa: a powerful pre-trained language model for Vietnamese. In Findings of the Association for Computational Linguistics: EACL 2023, A. Vlachos and I. Augenstein (Eds.), Dubrovnik, Croatia, pp. 1071–1078. External Links: Link, Document Cited by: §1, §2.2.
  • A. Vlachos and S. Riedel (2014) Fact checking: task definition and dataset construction. In Proceedings of the ACL 2014 Workshop on Language Technologies and Computational Social Science, Baltimore, MD, USA, pp. 18–22. External Links: Link, Document Cited by: §2.1.
  • T. Vu, D. Q. Nguyen, D. Q. Nguyen, M. Dras, and M. Johnson (2018) VnCoreNLP: a Vietnamese natural language processing toolkit. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Demonstrations, New Orleans, Louisiana, pp. 56–60. External Links: Link, Document Cited by: §3.5.3.
  • W. Y. Wang (2017) “liar, liar pants on fire”: a new benchmark dataset for fake news detection. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), Vancouver, Canada, pp. 422–426. External Links: Link, Document Cited by: §2.1.
  • Z. Wang, P. Ng, X. Ma, R. Nallapati, and B. Xiang (2019) Multi-passage bert: a globally normalized bert model for open-domain question answering. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pp. 5878–5882. Cited by: §2.2.
  • B. M. Yao, A. Shah, L. Sun, J. Cho, and L. Huang (2023) End-to-end multimodal fact-checking and explanation generation: a challenging dataset and models. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, pp. 2733–2743. Cited by: §4.2.3.
@@ Appendix A Data Generation Rules Analysis
Turn nouns into relative clauses E: Trong bốn nước xưa có nền văn minh lớn thì có ba nước xưa ở vào châu Á (Ấn Độ, Iraq (Lưỡng Hà) và Trung Quốc). (Among the four ancient civilizations, three were located in Asia (India, Iraq (Mesopotamia), and China).) C: Châu Á là nơi tồn tại ba nước có nền văn minh lớn. (Asia is the region where three major ancient civilizations existed.) 10.20%
Turn the object into relative clauses E: Lý thuyết "Quả cầu tuyết Trái Đất" cho rằng những sự thay đổi về mức độ CO2 vừa là nguyên nhân gây ra, vừa là nguyên nhân làm kết thúc thời kỳ cực lạnh ở cuối Liên đại Nguyên Sinh (Proterozoic). (The "Snowball Earth" theory posits that changes in the levels of CO2 were both the cause and the termination of the extreme cold period at the end of the Proterozoic Eon.) C: Theo lý thuyết "Quả cầu tuyết Trái Đất" cho thấy rằng sự thay đổi thành phần không khí mà cụ thể là lưu lượng CO2 là tác nhân cho kỷ băng giá, và cũng kết thúc luôn thời kỳ cực lạnh ở kỷ Proterozoic. (The "Snowball Earth" theory indicates that changes in the composition of the atmosphere, specifically the levels of CO2 which are the driving factors behind ice ages and the termination of extreme cold periods in the Proterozoic Eon.) 2.20%